OpenStudy (anonymous):

what is the wind chill with T=-10 cel and V=20m/sec.. using the wind formula? w=33-(10.45+10squV-V)(33-T)/22 @mathmale

OpenStudy (mathmale):

We're given the wind formula: \[ w=33-(10.45+10\sqrt{V} V-V)(33-T)/22\]. I'm not convinced this is correct. Please compare this to the original equation and let me know if my interpretation is incorrect.

OpenStudy (anonymous):

@mathmale the only thing that is wrong with the formula is that it should be one V under the \[\sqrt{v}\]

OpenStudy (mathmale):

So, is the following correct?\[w=33-(10.45+10\sqrt{V} -V)(33-T)/22?\] sorry this is taking so long, but until we're both sure we have the formula correct, there's no point in trying to evaluate it at some numeric value. If you were given this problem online, would you please take a screen shot of the original problem statement and share it with me?

OpenStudy (anonymous):

No it's okay.. thanks for the help.. the formula you just wrote is correct. T=-10 V=20 @mathmale

OpenStudy (mathmale):

OK: You are asked, "what is the wind chill with T=-10 cel and V=20m/sec.?" Just substitute these values of T and V into the equation.

OpenStudy (anonymous):

yes can you should me it worked out? because I think I did a step wrong.

OpenStudy (mathmale):

\[w=33-(10.45+10\sqrt{V} -V)(33-T)/22 ~when~T=-10 ~and~V=\frac{ 20 m }{ \sec }\]

OpenStudy (mathmale):

Let's evaluate this formula piece by piece. First, your (10.45 + 10 Sqrt(v) - V) becomes 10.45 + 10Sqrt(20m/sec) -20m/sec. This does not look right to me because the units of measurement are not consistent.

OpenStudy (mathmale):

(33-T) when T=-10 would be (33-[-10]) = 33+10=43.

OpenStudy (mathmale):

Again, I see inconsistencies in the units of measurement, and again encourage you to share a screen shot of the original problem.

OpenStudy (anonymous):

How do I screen shot it?

OpenStudy (mathmale):

First, in what material did you find this problem? If you're not sure how to take a screen shot, please just forward the URL of the online material in which you found the problem. How one takes a screen shot depends upon what computer and what software one has.

OpenStudy (anonymous):

4) The formula to find the wind chill temperature is given by Where, W is Wind Chill temperature (temperature with no wind) T is actual temperature in Celcius V is wind speeds in m/sec Find the Wind Chill temperature given the following: T =-10 , v = 20m/sec

OpenStudy (anonymous):

Just used the formula that you typed..

OpenStudy (mathmale):

Thanks. Still need to see the original formula. If you have a cell phone, consider taking a picture of the original formula and sharing the image with me. Alternatively, would YOU please substitute the given values into the formula, showing me your work? If necessary, use the Draw utility (below).

OpenStudy (mathmale):

I'm sorry, but as I explained earlier, I have my doubts that my interpretation of your formula is correct. Next week, in school, please ask a teacher or a friend to demonstrate for you how to take screen shots. Again, if you have a cell phone that can take photos, you have an alternative method for sharing images.

OpenStudy (shiraz14):

@mathmale: You may wish to refer to the following URL for evaluating the discrepancy between this formula and the original model: http://en.wikipedia.org/wiki/Wind_chill

OpenStudy (anonymous):

Yes that is the formula that @shiraz14 just shared with us from Wiki.. the orginial format @mathmale . Do you know how to do this @shiraz14 ?

OpenStudy (shiraz14):

@t.street17 : It's like what @mathmale has posted - you just need to substitute the values. @mathmale : I understand where you are coming from but I presume that we should just assume that they are dimensionless constants (to avoid confusion for now) ...

OpenStudy (mathmale):

@shiraaz14: It was VERY helpful of you to look up and share this formula. Thank you very much!

OpenStudy (shiraz14):

@mathmale : You're most welcome.

OpenStudy (mathmale):

Given that V=20m/sec, please evaluate \[(10\sqrt{V}-V+10.5).\]Please show your work.

OpenStudy (anonymous):

I just did that and I got 68.74 for the answer

OpenStudy (mathmale):

Next, given that T=-10 degrees Celcius, please evaluate\[33-T _{a}\]

OpenStudy (anonymous):

did you get the same answer I got @mathmale of 68.74?

OpenStudy (mathmale):

I haven't been calculating those myself, since I'm assuming you can do that. But in this case I'll check. Please more on to evaluating the other expression (see above).

OpenStudy (anonymous):

Yeah I have done all of that ... I am just checking my answer now

OpenStudy (mathmale):

My evaluation of \[(10\sqrt{V}-V+10.5)\] comes out to 35.22:

OpenStudy (anonymous):

mine was 35.17... let me check it

OpenStudy (mathmale):

\[10\sqrt{20}-20+10.5=44.72 - 20 + 10.5 = 24.72+10.5 = 35.22\]

OpenStudy (mathmale):

I'd say your 35.17 is close enough.

OpenStudy (mathmale):

Now, what about\[35-T _{a}?\]

OpenStudy (anonymous):

okay .... I think I am have the answer... this is making my head hurt... lol.. thanks for your help

OpenStudy (shiraz14):

@mathmale : I think there's a typo - it should be 33 - Ta?

OpenStudy (mathmale):

I'm so glad we've reached closure on this. Again, I'd suggest you ask around at school for help in taking screen shots, or experiment with taking and sharing images of your written work with your cell phone (assuming you have one). Sharing images would have made this conversation go a lot faster. Best of luck to you! And once again, thank you, @shiraz14!

OpenStudy (mathmale):

@shiraz14: again, you are correct. Thanks! I'm going to hire you as my consultant.

OpenStudy (shiraz14):

@mathmale : ;)

OpenStudy (shiraz14):

@t.street17 & @mathmale : There's only a slight variation between the model provided in the Wiki and the one provided by @t.street17 earlier. @t.street17, are you certain that this is the same formula provided in your worksheet? Or is the formula the same as the one in the Wiki?

OpenStudy (mathmale):

I noticed, also, that there are differences between t.street's formula and that provided in the Wiki. This is why I felt it so important to ask for an illustration of the problem you (@t.street) were working on. Your (@t.street's) formula includes division by 22 which is not a part of the Wiki-provided formula.

OpenStudy (shiraz14):

@mathmale : Yes, you're correct.

OpenStudy (shiraz14):

Goodbye for now